# A prime number “game of life”: can $\lfloor y\cdot p\#\rfloor$ be prime for all $p\geq 2$?

### v1

@@ 9. Broader utilization
$s$ $p$ $n$ $n*$ $s$ $p$ $n$ $n*$ $s$ $p$ $n$ $n*$
121 661 1701 105 156 911 5971 389 191 1153 19690 1147
122 673 1773 108 157 919 6366 404 192 1163 20364 1179
123 677 1801 116 158 929 6582 414 193 1171 21154 1216
124 683 1889 121 159 937 6799 433 194 1181 21954 1262
125 691 1885 127 160 941 7006 455 195 1187 22688 1287
126 701 1944 132 161 947 7151 463 196 1193 23528 1333
127 709 2008 138 162 953 7599 474 197 1201 24295 1363
128 719 2141 144 163 967 7920 494 198 1213 25085 1401
129 727 2136 150 164 971 8276 511 199 1217 25717 1442
130 733 2241 158 165 977 8708 524 200 1223 26680 1492
131 739 2354 165 166 983 9086 535 201 1229 27569 1527
132 743 2442 177 167 991 9577 553 202 1231 28242 1572
133 751 2548 179 168 997 9736 570 203 1237 28917 1614
134 757 2649 184 169 1009 10329 586 204 1249 29871 1659
135 761 2791 193 170 1013 10722 594 205 1259 30707 1708
136 769 2998 200 171 1019 11048 611 206 1277 31555 1754
137 773 3017 207 172 1021 11387 639 207 1279 32633 1808
138 787 3148 214 173 1031 11769 669 208 1283 33494 1871
139 797 3132 224 174 1033 12251 685 209 1289 34556 1921
140 809 3385 231 175 1039 12490 711 210 1291 35744 1981
141 811 3537 245 176 1049 12905 733 211 1297 36546 2034
142 821 3698 250 177 1051 13209 750 212 1301 37363 2087
143 823 3914 261 178 1061 13556 775 213 1303 38018 2150
144 827 4175 266 179 1063 13918 797 214 1307 38936 2218
145 829 4345 273 180 1069 14393 822 215 1319 39633 2303
146 839 4360 284 181 1087 15006 852 216 1321 40152 2375
147 853 4537 292 182 1091 15472 874 217 1327 40822 2456
148 857 4722 304 183 1093 15695 908 218 1361 42524 2541
149 859 4862 308 184 1097 16075 937 219 1367 44138 2626
150 863 4951 320 185 1103 16548 971 220 1373 45674 2704
151 877 5099 332 186 1109 16852 1008 221 1381 47308 2791
152 881 5228 345 187 1117 17548 1033 222 1399 49203 2876
153 883 5334 357 188 1123 17926 1064 223 1409 50999 2966
154 887 5460 364 189 1129 18415 1090 224 1423 53578 3047
155 907 5715 379 190 1151 19145 1124 225 1427 56009 3139
$s$ $p$ $n$ $n*$ $s$ $p$ $n$ $n*$ $s$ $p$ $n$ $n*$
226 1429 58245 3213 261 1663 141747 8687 296 1949 323493 35595
227 1433 60614 3340 262 1667 144129 8937 297 1951 332369 37696
228 1439 63012 3442 263 1669 146278 9203 298 1973 344077 40005
229 1447 65622 3520 264 1693 150011 9484 299 1979 355840 42551
230 1451 67609 3633 265 1697 153719 9799 300 1987 367371 45250
231 1453 69411 3747 266 1699 157438 10097 301 1993 379303 48293
232 1459 71677 3844 267 1709 160935 10460 302 1997 391133 51726
233 1471 73978 3960 268 1721 164713 10819 303 1999 402306 55407
234 1481 76816 4078 269 1723 168209 11187 304 2003 413306 59585
235 1483 78753 4173 270 1733 171914 11552 305 2011 424348 64450
236 1487 81051 4292 271 1741 175208 11966 306 2017 435084 69845
237 1489 82647 4411 272 1747 178924 12392 307 2027 446383 76167
238 1493 84106 4551 273 1753 182442 12861 308 2029 457067 83297
239 1499 85756 4684 274 1759 186286 13363 309 2039 468311 91883
240 1511 87469 4811 275 1777 190466 13889 310 2053 482068 102217
241 1523 90333 4958 276 1783 196294 14432 311 2063 494789 114639
242 1531 93284 5084 277 1787 200525 14980 312 2069 509095 130467
243 1543 96279 5233 278 1789 204944 15548 313 2081 524247 150449
244 1549 99150 5376 279 1801 209420 16163 314 2083 538695 177304
245 1553 101723 5539 280 1811 214672 16789 315 2087 551951 214976
246 1559 104560 5675 281 1823 218990 17483 316 2089 564045 271897
247 1567 107041 5832 282 1831 224272 18194 317 2099 577252 370885
248 1571 109795 5999 283 1847 230486 18982 318 2111 592642 592642
249 1579 111912 6155 284 1861 238339 19837
250 1583 114508 6341 285 1867 246127 20704
251 1597 117842 6528 286 1871 253236 21653
252 1601 121087 6716 287 1873 259946 22631
253 1607 124407 6901 288 1877 266707 23698
254 1609 126382 7113 289 1879 272114 24815
255 1613 128949 7305 290 1889 278512 26035
256 1619 130732 7493 291 1901 285821 27356
257 1621 132088 7720 292 1907 292121 28714
258 1627 133678 7965 293 1913 298404 30241
259 1637 135308 8188 294 1931 305779 31927
260 1657 137970 8439 295 1933 314060 33679
The calculation up to the point given above takes about five months with Pari/GP [10] on a single CPU core of a 2020 state-of-the-art PC with the following program, which in the given form starts at $s$ = 29 and keeps track of the numbers as a = $q_{s,1}$ and a memory friendly vector d comprising only the consecutive differences between the numbers of a given stage where d[1] = 0 and, for $i>1$, d[i] = $q_{s,i}-q_{s,i-1}$: 444 a and d can also be read in from a previously calculated output file
@@
Be sure to set allocatemem(108) or higher (and watch for the second “while” condition) when aiming for $s>318$.
@@
The ostensibly first titanic p#Y-prime that survives in the long run, the $16^{th}$ smallest of them all at $p=2357$, spanning its branch beyond 13147# ($s>1563$), is $q_{350,1}+$ 648890029246175571592953681722255270285895737015173251749453742478642415683875459 622820131990229693513724200589337798081947713889805053628824480433575300758724935 31089513889637480426037406091083077809690,
@@ 15. Acknowledgements
Many thanks to Charles R. Greathouse IV for his support when we first presented this sequence at Mersenneforum.org back in 2008. We are also deeply grateful to Alexei Kourbatov for helpful suggestions on the paper.

### v2

@@ 9. Broader utilization
$s$ $p$ $n$ nn* $s$ $p$ $n$ nn* $s$ $p$ $n$ nn*
121 661 1701 105 156 911 5971 387 191 1153 19690 1138
122 673 1773 108 157 919 6366 401 192 1163 20364 1169
123 677 1801 116 158 929 6582 411 193 1171 21154 1205
124 683 1889 121 159 937 6799 430 194 1181 21954 1251
125 691 1885 127 160 941 7006 452 195 1187 22688 1274
126 701 1944 132 161 947 7151 460 196 1193 23528 1319
127 709 2008 138 162 953 7599 471 197 1201 24295 1349
128 719 2141 144 163 967 7920 491 198 1213 25085 1387
129 727 2136 150 164 971 8276 508 199 1217 25717 1428
130 733 2241 158 165 977 8708 521 200 1223 26680 1477
131 739 2354 165 166 983 9086 532 201 1229 27569 1510
132 743 2442 177 167 991 9577 550 202 1231 28242 1555
133 751 2548 179 168 997 9736 566 203 1237 28917 1595
134 757 2649 184 169 1009 10329 581 204 1249 29871 1639
135 761 2791 193 170 1013 10722 589 205 1259 30707 1688
136 769 2998 200 171 1019 11048 606 206 1277 31555 1733
137 773 3017 207 172 1021 11387 634 207 1279 32633 1786
138 787 3148 214 173 1031 11769 663 208 1283 33494 1849
139 797 3132 224 174 1033 12251 679 209 1289 34556 1899
140 809 3385 231 175 1039 12490 705 210 1291 35744 1957
141 811 3537 245 176 1049 12905 727 211 1297 36546 2010
142 821 3698 250 177 1051 13209 743 212 1301 37363 2063
143 823 3914 261 178 1061 13556 768 213 1303 38018 2123
144 827 4175 266 179 1063 13918 790 214 1307 38936 2188
145 829 4345 273 180 1069 14393 815 215 1319 39633 2272
146 839 4360 284 181 1087 15006 845 216 1321 40152 2344
147 853 4537 291 182 1091 15472 867 217 1327 40822 2420
148 857 4722 303 183 1093 15695 901 218 1361 42524 2500
149 859 4862 307 184 1097 16075 930 219 1367 44138 2580
150 863 4951 319 185 1103 16548 963 220 1373 45674 2654
151 877 5099 331 186 1109 16852 999 221 1381 47308 2738
152 881 5228 344 187 1117 17548 1024 222 1399 49203 2818
153 883 5334 356 188 1123 17926 1055 223 1409 50999 2905
154 887 5460 363 189 1129 18415 1081 224 1423 53578 2982
155 907 5715 377 190 1151 19145 1115 225 1427 56009 3071
$s$ $p$ $n$ nn* $s$ $p$ $n$ nn* $s$ $p$ $n$ nn*
226 1429 58245 3143 261 1663 141747 8077 296 1949 323493 26835
227 1433 60614 3264 262 1667 144129 8290 297 1951 332369 28016
228 1439 63012 3359 263 1669 146278 8514 298 1973 344077 29288
229 1447 65622 3434 264 1693 150011 8752 299 1979 355840 30683
230 1451 67609 3542 265 1697 153719 9016 300 1987 367371 32050
231 1453 69411 3650 266 1699 157438 9268 301 1993 379303 33504
232 1459 71677 3745 267 1709 160935 9582 302 1997 391133 35062
233 1471 73978 3854 268 1721 164713 9877 303 1999 402306 36665
234 1481 76816 3964 269 1723 168209 10187 304 2003 413306 38433
235 1483 78753 4049 270 1733 171914 10497 305 2011 424348 40351
236 1487 81051 4156 271 1741 175208 10838 306 2017 435084 42411
237 1489 82647 4260 272 1747 178924 11177 307 2027 446383 44703
238 1493 84106 4386 273 1753 182442 11549 308 2029 457067 47035
239 1499 85756 4511 274 1759 186286 11951 309 2039 468311 49593
240 1511 87469 4631 275 1777 190466 12354 310 2053 482068 52470
241 1523 90333 4765 276 1783 196294 12773 311 2063 494789 55576
242 1531 93284 4879 277 1787 200525 13184 312 2069 509095 58953
243 1543 96279 5012 278 1789 204944 13625 313 2081 524247 62650
244 1549 99150 5141 279 1801 209420 14097 314 2083 538695 66722
245 1553 101723 5281 280 1811 214672 14574 315 2087 551951 71192
246 1559 104560 5402 281 1823 218990 15100 316 2089 564045 76342
247 1567 107041 5542 282 1831 224272 15623 317 2099 577252 81969
248 1571 109795 5691 283 1847 230486 16200 318 2111 592642 88171
249 1579 111912 5827 284 1861 238339 16792 319 2113 605293 95404
250 1583 114508 5989 285 1867 246127 17405 320 2129 620280 103655
251 1597 117842 6160 286 1871 253236 18037 321 2131 635284 113441
252 1601 121087 6327 287 1873 259946 18726 322 2137 649598 124938
253 1607 124407 6493 288 1877 266707 19481 323 2141 664112 138630
254 1609 126382 6682 289 1879 272114 20242 324 2143 676043 155455
255 1613 128949 6854 290 1889 278512 21025 325 2153 689579 176551
256 1619 130732 7021 291 1901 285821 21851 326 2161 703065 203256
257 1621 132088 7226 292 1907 292121 22714 327 2179 720414 240371
258 1627 133678 7440 293 1913 298404 23666 328 2203 742364 292120
259 1637 135308 7642 294 1931 305779 24693 329 2207 763990 370783
260 1657 137970 7862 295 1933 314060 25728 330 2213 787038 506979
Data for $s=331$ as output from the program on the following page is given as an ancillary file to this version of the paper.
[…]
@@
[…]
@@
The ostensibly first titanic p#Y-prime that survives in the long run, the $16^{th}$ smallest of them all at $p=2357$, spanning its branch beyond 13751# ($s>1626$), is $q_{350,1}+$ 648890029246175571592953681722255270285895737015173251749453742478642415683875459 622820131990229693513724200589337798081947713889805053628824480433575300758724935 31089513889637480426037406091083077809690,
@@ 15. Acknowledgements
[…]